OpenStudy (anonymous):

Identify the vertical asymptotes of f(x) = 10 over quantity x squared minus 7x minus 30. x = 10 and x = 3 x = 10 and x = -3 x = -10 and x = 3 x = -10 and x = -3

OpenStudy (anonymous):

\[f(x) = \frac{ 10 }{ x^2 -7x -30 }\]

OpenStudy (misty1212):

set \[x^2-7x-30=0\] solve for \(x\)

OpenStudy (anonymous):

okay give me a sec

OpenStudy (anonymous):

fortunalely for you this factors as \[(x-10)(x+3)=0\] so the zeros are easy to find

OpenStudy (misty1212):

yeah @satellite73 gave you a huge hint there

OpenStudy (anonymous):

so its x = -10 and x = 3

OpenStudy (misty1212):

no dear

OpenStudy (misty1212):

\[x-10=0\\ x=10\]

OpenStudy (misty1212):

also \[x+3=0\\ x=-3\]

OpenStudy (anonymous):

oh so i have to bring it to the other side okay i see

OpenStudy (misty1212):

riiiiiigght

OpenStudy (anonymous):

thnx i have one more do you mind helping with?

OpenStudy (misty1212):

sure

OpenStudy (anonymous):

Identify the horizontal asymptote of f(x) = quantity 2 x minus 1 over quantity x squared minus 7 x plus 3. y = 0 y = 1 over 2 y = 2 No horizontal asymptote

OpenStudy (anonymous):

\[f(x) = \frac{ 2x - 1 }{ x^2 - 7x +3 }\]

OpenStudy (anonymous):

do i basically do the same thing as before?

OpenStudy (misty1212):

no!!

OpenStudy (misty1212):

horizontal not vertical

OpenStudy (anonymous):

oh so what do i need to do?

OpenStudy (misty1212):

the degree of the numerator is less than the degree of the denominator

OpenStudy (misty1212):

the degree of the numerator is 1, the degree of the denominator i s2

OpenStudy (misty1212):

that means the denominator grows faster than the numerator so the horizontal asymptote is \[y=0\]

OpenStudy (anonymous):

thank you so much!!!

OpenStudy (anonymous):

i gtg but i have lie 5 more questions to do later can you help with?

OpenStudy (misty1212):

sure dear if i am on